OpenStudy (anonymous):

If A=(1,2,3) and B=(4,5), find A x B and B x A. Determine n(A x B) and n(B x A). Could you please explain to me how to do these type of questions !

OpenStudy (anonymous):

are these matrices???

OpenStudy (anonymous):

these are sets

OpenStudy (anonymous):

B x A=A x B = 1x4 +1x5 +2x4 + 2x5 + 3x4 + 3x5 n(A x B) = n x A x B

OpenStudy (amistre64):



OpenStudy (amistre64):

BxA is the set of ordered points like normal now; (b,a)

OpenStudy (amistre64):

X x Y is what your used to seeing as a graph; with all the ordered pairs (x,y)

OpenStudy (zarkon):

\[A\times B=\{(a,b)|a\in A, b\in B\}\] \[A\times B=\{(1,4),(2,4),(3,4),(1,5),(2,5),(3,5)\}\]